A 2.0 mol. sample of an ideal gas is kept at 1.0â—¦c during an expansion from 2.0 L to 9.0 how much energy transfer by heat occurs with the surroundings in this process?


1
Expert's answer
2021-05-04T06:29:08-0400

Answer :-


give

mole = 2

VA= 2 L

VB = 9 L

T = 1oC

R = 8.3 J/K

first law of thermodynamics is 

"\u0394\n\n\n\nU=Q\u2212W"


 if the temperature is held constant, the internal energy of the system also is constant, and so 

ΔU=0

. From the first law of thermodynamics, it follows that 

"\boxed{Q=−W}"

 for this same isothermal process.


also work

"\\boxed{W_{A\u2192B}=nRTln\\frac{\u2061V_B}{V_A }}"



"Q = - nRTln\\frac{\u2061V_B}{V_A }"


="(2)\\times(8.3)\\times(273)ln\\frac{\u20619}{2 }"


= 6797.7 joule